OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = -4(x + 6)2 + 2?

terenzreignz (terenzreignz):

Can't get much better than this :) If your quadratic equation is of the form... \[\LARGE y = a(x-\color{red}h)^2+\color{blue}k\] Then your vertex is the point... \[\LARGE (\color{red}h,\color{blue}k)\]

OpenStudy (anonymous):

i still dont understand

terenzreignz (terenzreignz):

In that form of the equation (which, understandably, is called the vertex form), we have... \[\LARGE y = -4(x+6)^2+2\] In this scenario, which numbers are our h and k?

terenzreignz (terenzreignz):

\[\LARGE y = a(x-\color{red}h)^2+\color{blue}k\]

OpenStudy (anonymous):

6=h 2=k

terenzreignz (terenzreignz):

Not quite, but close... as you can see, in the standard form, h is positive if it is accompanied by a minus (-) sign... so if in your equation, 6 is accompanied by a plus (+) sign, then h must be equal to...?

OpenStudy (anonymous):

positive 6

terenzreignz (terenzreignz):

Nope. if h = +6, then our equation has the form... \[\LARGE y = a(x-\color{red}h)^2+\color{blue}k\] \[\LARGE y = a(x-\color{red}{(+6)})^2+\color{blue}2\] \[\LARGE y = a(x-\color{red}6)^2+\color{blue}2\quad \ne\quad a(x+6)^2+2\]